OpenStudy (anonymous):

Trigonometric behaviors- Tangents:

OpenStudy (shayaan_mustafa):

sure friend.

OpenStudy (anonymous):

Btw, its only question1 not all of them:)

OpenStudy (shayaan_mustafa):

let us solve Q#1

OpenStudy (anonymous):

Yes:)

OpenStudy (shayaan_mustafa):

Write the given question in the general form as.\[\large y=Atan[B(x-(\frac{C}{B}))]+C\]\[\large y=-2\tan[3(x-(-\frac{\pi}{3}))]+3\] Now compare results amplitude=|A|=-2 Period=2*pi/B 2*pi/3

OpenStudy (shayaan_mustafa):

Thnx for testimonial friend. :)

OpenStudy (anonymous):

Wait, but how come there is a pi/3 in the equation?

OpenStudy (shayaan_mustafa):

it don't have vertical asymptote but two horizontal asymptote.

OpenStudy (shayaan_mustafa):

180degree = pi radians.

OpenStudy (anonymous):

Oh thats right:)

OpenStudy (anonymous):

so the period is 2*pi/3

OpenStudy (shayaan_mustafa):

yes.

OpenStudy (anonymous):

Ok. Is this range correct? {y | y is an element or R}

OpenStudy (shayaan_mustafa):

yes. range will be all real numbers.

OpenStudy (anonymous):

What would my domain be? This one in particular I am having the most trouble with

OpenStudy (campbell_st):

one simple problem is that tan has a period of pi... and not 2pi... so the period is pi/3 for the tan graph... the -2 means there curve is reversed +3 means the centre of the curve is moved up to the line y = 3